OpenStudy (anonymous):

Evaluate the polynomial 6x - 7 for 1/2n^2 - 1/2n and y=2

OpenStudy (accessdenied):

um... is there missing information? There's no relation between x, n, and y that I know of... D:

OpenStudy (anonymous):

i was wondering the same thing but that is all it gives me lol the answer choices are A. 15 B. –20 C. 4 D. –16

OpenStudy (precal):

surely there is a typo :)

OpenStudy (anonymous):

no this is the problem

OpenStudy (precal):

You might want to double check. When you evaluate, you usually have the same variables. example evaluate the following expression 2ab +3 a=2 b=4

OpenStudy (accessdenied):

i tried to equate the polynomial back to each of the answers, none of the x-values stick out to me as particularly helpful 6x - 7 = 15 6x = 22 x = 22/6 = 11/3 6x - 7 = -20 6x = -13 x = -13/6 6x - 7 = 4 6x = 11 x = 11/6 6x - 7 = -16 6x = -9 x = -9/6 = -3/2

OpenStudy (anonymous):

i failed my paper so no i have to retake it

OpenStudy (anonymous):

THE ANSWER IS -20 JUST CALLED MY SCHOOL ITS A TYPO!!!!!!!!!!!!!!!!!